OpenStudy (anonymous):

Solve the equation2x^2-3x-5√(2x^2-3x+2)+6=0 . . . ANSWER = {1,1/2,-2,7/2}

OpenStudy (anonymous):

anybody there

OpenStudy (anonymous):

\[equation:- 2x^2-3x-5\sqrt{2x^2-3x+2}+6=0\]

OpenStudy (anonymous):

\[2x^2-3x-5\sqrt{2x^2-3x+2}+6=0\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

then how will i get the answer

OpenStudy (anonymous):

\[2 x^2-3 x+6-5 \sqrt{2 x^2-3 x+2}=0 \]\[2 x^2-3 x+6=5 \sqrt{2 x^2-3 x+2} \]\[(2 x^2-3 x+6)^2=(5 Sqrt[2 x^2-3 x+2])^2 \]\[4 x^4-12 x^3-17 x^2+39 x-14=0 \]\[(x-1) (x+2) (2 x-7) (2 x-1)=0 \]

OpenStudy (anonymous):

The Book says that the answer is {1,1/2,-2,7/2}

OpenStudy (anonymous):

hello anybody there? help plz

OpenStudy (anonymous):

If x-1=0, then x=1 If x+2=0 then x=-2 If 2x-7=0 then x= 7/2 if 2x-1=0 then x=1/2